OpenStudy (anonymous):

A person takes a trip, driving with a constant speed 89 km/h except for a 16.4 min rest stop. If the person’s average speed is 75.1 km/h, how much time is spent on the trip? Answer in units of h

OpenStudy (kropot72):

Let the distance travelled be d and the total time taken be t hours. Average speed = d/t Therefore d = 75.1t ...............................(1) d = 89(t - 16.4/60) ................................(2) Substituting the value for d in equation (1) into equation (2) gives 75.1t = 89(t - 16.4/60) ..........................(3) Now you need to solve equation (3) to find the value of t.